# A Survey of Protocol Fuzzing

### v2

@@ 4.3. Testcase Construction
The construction strategy used in the protocol fuzzing can be categorized into packet-level and sequence-level.
@@
4.3.1. Packet-Level Construction Strategy (labeled as “Packet” in Table 3 column 6)
@@
Packet-Level construction strategies of protocol fuzzers basically inherit the strategies of general fuzzers, such as bit flip, set zero, etc. In this paragraph, more consideration is given to the construction strategies that leverage protocol-specific characteristics for reducing input space or improving the effectiveness of triggering bugs. For the former purpose (i.e., reducing input space), SPIDER(Li et al., 2022b) leverages the domain-specific insight that most of the Openflow messages trigger new system events in existing SDN controllers that affect state computation and resource footprints. Based on the insight, SPIDER can directly generate event sequences rather than generating Openflow messages, which significantly narrows down the input space. Also, L2Fuzz (Park et al., 2022) divides the L2CAP packet format into the field that can be mutated and keeps the other fields unchanged to generate testcases that are less likely to be rejected. IPSpex (Zheng and Sun, 2022) combines network traffic and execution traces of network packet construction to extract ICS protocol message field semantics. The strategies targeting the later purpose (i.e., improving the effectiveness of triggering bugs) are mostly heuristics summarized from practices. For example, EmNetTest (Amusuo et al., 2023) systematically generates validly constructed packets with invalid header fields or truncated headers. The insight behind this strategy is gained from a comprehensive study of 61 reported vulnerabilities in Embedded Network Stacks (ENS). Similar strategies are mentioned in many industry conference works. BadMesher (Qu et al., 2021) adopts several domain-specific strategies such as setting the length field to margin values, and randomly deleting some fields, to improve the effectiveness of triggering bugs in Wi-Fi mesh devices. Yen et al. (Yen et al., 2021) find that some strategies such as mutating the ID field to a non-existing ID, changing to port number or length field to a boundary value (e.g., 0xFF/0x00), and changing IP to some random addresses, can be quite effective in fuzzing Data Distribution Service (DDS) protocol. Similarly, BrokenMesh (Yan et al., 2022) adopts some strategies like mutating the packet count or the length field in fuzzing the Bluetooth Mesh protocol.
@@
4.3.2. Sequence-Level Construction Strategy (labeled as “Sequence” in Table 3 column 6)
@@
Generation-Based Fuzzers: These fuzzers construct message sequences leveraging established protocol knowledge, such as standard state machines and inter-message dependency relationships. Notable examples include works (Beurdouche et al., 2015; Somorovsky, 2016; Vanhoef, 2017; Reen and Rossow, 2020) that generate aberrant message traces by applying strategies like the addition or removal of random protocol messages to valid sequences derived from standard state machines. Projects like Sweyntooth (Garbelini et al., 2020), Greyhound (Garbelini et al., 2022c), and Braktooth (Garbelini et al., 2022a) meticulously monitor state transitions of PUT and strategically inject valid packets at incorrect states to elicit anomalies, in accordance with the state machine model. Recent research by Fiterau-Brostean et al.(Fiterau-Brostean et al., 2023) proposes a novel method for detecting state machine bugs by inputting a catalogue of finite automatons which indicate certain types of state machine bugs, as well as a model of the PUT’s. It can then analyze the models and produces testcases that expose the bug.
Mutation-Based Fuzzers: These fuzzers predominantly employ simple yet effective strategies to mutate the message sequences of seeds. This includes techniques such as packet shuffling, random insertion, or deletion. For instance, AFLNET (Pham et al., 2020) constructs message sequences by maintaining a pool of messages from network traces that can be integrated into or substituted for existing seeds. AFLNET further employs a blend of byte-level and sequence-level operators, including replacement, insertion, duplication, and deletion of messages, to craft the message sequences. Similarly, DYFuzzing (Amusuo et al., 2023) mutates seeds and applies Dolev-Yao (DY) attacker strategies. Frankenstein (Ruge et al., 2020) reorganizes known message sequences to enhance code coverage. He et al. (He et al., 2022) propose a unique fuzzer for the 5G non-access-stratum (NAS) protocol, which extracts packets from captures into a structured message table. This fuzzer then applies differentiated mutation strategies to various key fields, thereby significantly enhancing the intelligence and precision of the message mutation process. It is important to note that mutation-based fuzzers must judiciously manage the correlation of specific fields in message sequences, such as session numbers, counters, or timestamps. Indiscriminate mutations in these fields could render the input ineffective and lead to early rejection. To address this challenge, AFLNET (Pham et al., 2020) modifies the code of the PUT to use a fixed session number, thereby ensuring the effectiveness of the fuzzing process.
@@ 5. Executor
| Year | Work | T | Target | Efficient Execution | Runtime Info Extraction | ||||||
@@
| 2020 | Frankenstein(Ruge et al., 2020) | ◐ | Bluetooth | SE | Shared memory | VMSR | - | Code Cov | SDI | ||
@@
| 2020 | ToothPicker(Heinze et al., 2020) | ◐ | Bluetooth | HIL | FHPI | ThrdR | - | Code Cov | SDI | ||
| 2021 | ICSFuzzer(Fang et al., 2021) | ● | ICS | - | Socket (P2P) | ProcR | - | Exec Traces | SDI | ||
@@
| 2021 | TCP-Fuzz(Zou et al., 2021) | ◐ | TCP | - | Socket (P2P) | - | - | Branch Cov | SSI | ||
@@
○: Whitebox Fuzzer; ●: Blackbox Fuzzer; ◐: Greybox Fuzzer; T: Taxonomy; HIL: Hardware-In-the-Loop; General: The fuzzer is not designed for a specific type of protocol; CUVM: Commonly Used Virtual Machine; SE: Specialized Emulation; OTA: Over-the-air; UDS: Unix Domain Socket; MiTM: Man-in-The-Middle-based packet injection; VMSR: Virtual Machine-level Snapshot and Recovery mechanism; ProcR: Process Restart; PhyR: Physical Reset; ThrdR: Thread Restart; DR: Database Reset; PSR: Process-level Snapshot and Recovery mechanism; UPSR: User-Provided Script Reset; HA: Hardware-Assisted mechanism; EOB: Externally-Observable-Behavior-based method; SDI: Software Dynamic Instrumentation; SSI: Software Static Instrumentation; BAF: Bandwidth Amplification Factor; -: Not implemented; RM: Reset Message; IMFR: In-Memory Filesystem Reset; Resp: Responses.
In this section, we will introduce the key improvements of protocol fuzzers on the executor in detail. As shown in Fig. 6, an executor in protocol fuzzing normally includes four key processes. First, the executor needs to prepare an executable execution environment for PUT (①. Execution Environment Preparation), and then send input to PUT through the input feeding mechanism (②. Input Feeding), extract runtime information during the input processing (③. Information Extraction), and reset the execution state and environment state to a specific state after the execution of the current iteration is completed (④. Execution Reset).
@@ 5.1. Efficient Execution
In protocol fuzzing, there are commonly two directions to improve the fuzzing efficiency: 1) establishing an execution environment that enables parallel testing (① in Fig. 6); 2) minimizing the execution cost of each iteration of testing (② and ④ in Fig. 6).
@@ 5.1.1. Scalability Improvement
Scalable fuzzing, in this context, refers to the capacity to create multiple testing environments for parallel fuzzing. This is crucial in protocol fuzzing, where many fuzzing targets are closely bounded to hardware. The traditional approach for parallel testing of purchasing multiple physical devices can be economically burdensome and inefficient. For protocol fuzzing, since many fuzzing targets depend on specialized execution environments, concurrent testing of these targets can only be carried out by purchasing multiple physical devices, which leads to high economic costs and waste.
@@
However, two difficulties hinder the usage of emulation in protocol fuzzing. The first is the availability of the protocol implementation binary as many firmware images are not publicly available. Second, compared to the diversity of hardware, existing emulators can only support a small fraction of them. These difficulties lead to a lot of work still performing fuzzing in a hardware-in-the-loop way (labeled as “HIL” in Table. 5 column 4) (Luo et al., 2019; Garbelini et al., 2020; Feng et al., 2021; Garbelini et al., 2022c; Gorenc and Molinyawe, 2014; Park et al., 2022).
There are also some works addressing these issues according to the characteristics of different devices (labeled as “SE” in Table. 5 column 4). For the first challenge, the existing works obtain the target binaries by intercepting Over-The-Air (OTA) firmware updates, or extracting using vendor-specific command or debugging ports. For example, Frankenstein (Ruge et al., 2020) leverages the Patchram mechanism, a Broadcom vendor-specific command that can be used to temporarily patch breakpoints to the ROM, to take the memory snapshot of a physical Bluetooth chip and emulate it in an unmodified version of QEMU. To address the second question, the existing work often uses an approach called rehosting to partially emulate the functionality of the physical hardware. For example, BaseSafe (Maier et al., 2020) selectively rehosts several parser functions of signaling messages leveraging Unicorn engine, which is a popular CPU emulator (Quynh and Vu, 2015).
@@ 5.1.2. Execution Cost Reduction
Another direction to improve fuzzing efficiency is to optimize the intermediate execution steps in each iteration. Below we present the progress of existing works toward this direction, which mainly focus on three sub-procedures: input feeding (② in Fig. 6) and execution reset (④ in Fig. 6), specifically.
@@
Input Feeding. Input feeding mechanism acts as a pipeline between the input generator and the PUT to pass the testcase to the PUT for parsing and execution. According to the Inter-Process Communication (IPC) mechanisms that the communication between Fuzzer and PUT rely on, the existing approaches can be roughly divided into four categories, namely OTA-based, socket-based, shared-memory-based, and file-based approaches. OTA- and socket-based approaches are mostly used when the fuzzer and PUT cannot be deployed on the same physical device. The latter two approaches can be used to speed up input feeding when PUT and the fuzzer can be deployed on the same device.
@@
OTA-Based Input Feeding (labeled as “OTA” in Table. 5 column 5). In general, OTA-based input feeding mechanisms are mostly used in the scenario of fuzzing the implementations of protocols that are typically closed in nature and tightly integrated with hardware components, such as Wi-Fi (Garbelini et al., 2022c; Vanhoef, 2017; Sommer et al., 2018), Bluetooth (including classic Bluetooth and BLE) (Garbelini et al., 2022a; Park et al., 2022; Garbelini et al., 2020), LTE (Maier et al., 2020), 4G/5G (Garbelini et al., 2022b; He et al., 2022) and SMS/MMS protocols (Gorenc and Molinyawe, 2014). In this approach, PUT and the fuzzer need to be deployed in adjacent physical spaces and communicate with each other on specific frequency bands. Thus, OTA-based fuzzers require the use of radio-frequency transceiver devices with receive and transmit functions, such as a Software-Defined Radio (SDR), to handle signals over a wide tuning range. OTA-based fuzzing provides the capability of testing the entire protocol stack including the physical layer. However, OTA-based approaches are the slowest among the above-mentioned approaches. Therefore, many wireless protocol fuzzers try to use other input feeding mechanisms to have a better performance, which will be introduced in the following.
@@
Socket-Based Input Feeding (labeled as “Socket” in Table. 5 column 5). Socket-based input feeding mechanisms are mostly used in protocol implementations based on TCP/IP infrastructures. In common cases under these approaches, fuzzer and PUT communicate with each other through IP addresses, via socket mechanisms including TCP socket and UDP socket. The socket-based approaches include two deployment modes, one is point-to-point (P2P) communication between the fuzzer and PUT (Chen et al., 2018; Walz and Sikora, 2017; Fiterau-Brostean et al., 2020; Pham et al., 2020; Natella, 2022; Zou et al., 2021; Meng et al., 2021; Li et al., 2022a; Andronidis and Cadar, 2022). The fuzzer can play the role of a client or server depending on the role of the PUT. The other deployment mode is Man-in-the-middle (MiTM), in which the fuzzer acts as a proxy between two communication parties and performs mutation or injection to the normal communication traffic (Jero et al., 2018; Peterson et al., 2020). The MiTM-based input feeding is mostly used in the scenario where the protocol involves certain contextual information (checksum, packet sequences, etc.) that cannot keep valid by mutating static seeds. However, both modes need to address two challenges. First, socket communication is quite heavy and involves lots of context switches. Existing works improve the efficiency of the socket-based input feeding mechanisms by avoiding the use of these expensive network functions. For example, SnapFuzz (Andronidis and Cadar, 2022) replaces the original internet socket with UNIX domain socket (Coffield and Shepherd, 1987), a lightweight IPC mechanism that does not have the routing, checksum calculation operations that IP sockets have. Second, it’s hard for the fuzzer to determine whether the PUT has already finished processing the previous message and is ready to receive the next message. The PUT may reject the messages coming too early when the target is not ready, thus causing the fuzzer to desynchronize from its state machine. To solve this issue, Fiterau-Brostean et al. (Fiterau-Brostean et al., 2020) and AFLNET (Pham et al., 2020) set static time intervals to wait for the PUT to initialize, process requests, and send responses. However, static timers are too coarse-grained and can waste a lot of time waiting for the timeout, thus slowing down the fuzzing process. SnapFuzz (Andronidis and Cadar, 2022) and AMPFuzz (Krupp et al., 2022) develop a more fine-grained method to inspect the state of the socket. Specifically, they use the function call to related network system calls such as , as a sign of ready to receive the next message. They monitor all these function calls through binary rewriting and compile-time code instrumentation, and then notify the fuzzer to send the next iteration of input.
@@
File-Based Input Feeding (labeled as “File” in Table. 5 column 5). File-based input feeding leverages static or dynamic instrumentation techniques to replace heavy network operations with file operations to achieve a performance boost. For example, Yurong et al. (Chen et al., 2019) transform socket communication to file operations using preloading customized libraries (zardus, 2019) under the circumstance that the source code of target is not available. Similarly, Nyx-net (Schumilo et al., 2022) injects a library into the target to hook the network functions of the target connection to obtain their associated file descriptors and injected fuzzing input to the right place.
@@
Shared-Memory-Based Input Feeding (labeled as “Shared-Memory” in Table. 5 column 5). Shared-Memory-Based input feeding writes the fuzzing input to the address of shared memory and hooks the related functions to read the testcase from shared memory (Maier et al., 2020; Ruge et al., 2020; Fu et al., 2023; Bars et al., 2023). For example, BaseSafe (Maier et al., 2020) executes each generated testcase in a forked copy of the target process, and the input for each run is copied to the appropriate address in the corresponding child process. Similarly, Frankenstein (Ruge et al., 2020) creates a virtual modem to inject custom packets. The fuzzed input are written to the receive buffer in RAM that is mapped to the hardware receive buffer using direct memory access (DMA). Also, HNPFuzzer (Fu et al., 2023) emulate network functions based on shared memory to short the time consumption due to message transmission between fuzzer and PUT.
@@
Others. There are also works that rely on specialized communication channels to feed fuzzing inputs. For example, to fuzz the client of the Remote Desktop Protocol (RDP), Park et al. leverage the virtual channel, an abstraction layer in RDP that is used for transporting data, to actively send fuzzing input from the server to the client (Park et al., 2019). Song et al. use a media converter to convert the traffic between Automotive Ethernet and standard Gigabit Ethernet and fuzz the SOME/IP protocol stack of the electronic control unit (ECU) (Song et al., 2022), which is a control communication protocol between ECUs.
@@
The reset of execution includes three key steps, which are reset time selection, execution state reset and execution environment reset. First of all, the executor needs to judge whether the current iteration is over (1. reset time selection) before execution reset. After confirming that the current iteration of execution has finished, the executor needs to separately reset the runtime state of PUT (2. execution state reset) and its impact on the external execution environment such as the file system and database (3. execution environment reset). Below we summarize the progress of existing works in these three key steps separately.

### v3

@@ 4.3. Testcase Construction
The construction strategy used in protocol fuzzing can be categorized into packet-level and sequence-level.
@@
4.3.1. Packet-Level Construction Strategy (labeled as “Packet” in Table LABEL:table-inputgen column 6)
@@
Packet-Level construction strategies of protocol fuzzers basically inherit the common strategies of general fuzzers. For example, elements such as relation, fixup and transform in Peach Fuzzer (Eddington, 2014) are often used to describe the relationships between protocol fields such as length, checksums, and encoding transformations. General mutation methods like bit flip and set zero are also commonly used in generating packet-level testcases. In this paragraph, more consideration is given to the construction strategies that take advantage of protocol-specific characteristics to reduce input space or improve the effectiveness of bug triggering. For the former purpose (i.e., reducing input space), SPIDER(Li et al., 2022b) leverages the domain-specific insight that most Openflow messages trigger new system events in existing SDN controllers that affect state computation and resource footprints. Based on the insight, SPIDER can directly generate event sequences rather than generating Openflow messages, significantly reducing the input space. In addition, L2Fuzz (Park et al., 2022) divides the L2CAP packet format into the field that can be mutated and keeps the other fields unchanged to generate testcases that are less likely to be rejected. IPSpex (Zheng and Sun, 2022) combines network traffic and execution traces of network packet construction to extract the message field semantics of ICS protocols. Strategies targeting the later purpose (i.e., improving the effectiveness of triggering bugs) are mainly heuristics summarized from practices. For example, EmNetTest (Amusuo et al., 2023) systematically generates validly constructed packets with invalid header fields or truncated headers. The insight behind this strategy is gained from a comprehensive study of 61 reported vulnerabilities in Embedded Network Stacks (ENS). Similar strategies are mentioned in many industry conference works. BadMesher (Qu et al., 2021) adopts several domain-specific strategies, such as setting the length field to margin values, and randomly deleting some fields, to improve the effectiveness of triggering bugs in Wi-Fi mesh devices. Yen et al. (Yen et al., 2021) find that some strategies such as mutating the ID field to a nonexisting ID, changing the port number or length field to a boundary value (e.g., 0xFF/0x00), and changing IP to some random addresses, can be quite effective in fuzzing Data Distribution Service (DDS) protocol. Similarly, BrokenMesh (Yan et al., 2022) adopts some strategies such as mutating the packet count or the length field in fuzzing the Bluetooth Mesh protocol. TaintBFuzz (Ren et al., 2023) employs static taint analysis to identify the connections between Zigbee protocol fields and the variables used to make path decisions in Zigbee implementations, thereby prioritizing mutations that are anticipated to reveal additional code paths and improve the efficacy of testing. MPFuzz (Luo et al., 2024b) uses a global synchronization mechanism to share key field information across different fuzzing instances and performs targeted mutations based on the semantic characteristics of these fields, making it easier to discover potential vulnerabilities.
@@
4.3.2. Sequence-Level Construction Strategy (labeled as “Sequence” in Table LABEL:table-inputgen column 6)
@@
Generation-Based Fuzzers: These fuzzers construct message sequences leveraging established protocol knowledge, such as standard state machines and inter-message dependency relationships. Notable examples include works (Beurdouche et al., 2015; Somorovsky, 2016; Vanhoef, 2017; Reen and Rossow, 2020) that generate aberrant message traces by applying strategies like the addition or removal of random protocol messages to valid sequences derived from standard state machines. Projects like Sweyntooth (Garbelini et al., 2020), Greyhound (Garbelini et al., 2022c), and Braktooth (Garbelini et al., 2022a) meticulously monitor state transitions of PUT and strategically inject valid packets at incorrect states to elicit anomalies, in accordance with the state machine model. Recent research by Fiterau-Brostean et al.(Fiterau-Brostean et al., 2023) proposes a novel method for detecting state machine bugs by inputting a catalog of finite automatons which indicate certain types of state machine bugs, as well as a model of the PUT’s. It can then analyze the models and produces testcases that expose the bug.
Mutation-Based Fuzzers: These fuzzers predominantly employ simple yet effective strategies to mutate the message sequences of seeds. This includes techniques such as packet shuffling, random insertion, or deletion. For instance, AFLNET (Pham et al., 2020) constructs message sequences by maintaining a pool of messages from network traces that can be integrated into or substituted for existing seeds. AFLNET further employs a blend of byte-level and sequence-level operators, including replacement, insertion, duplication, and deletion of messages, to craft the message sequences. Similarly, DYFuzzing (Amusuo et al., 2023) mutates seeds and applies Dolev-Yao (DY) attacker strategies. Frankenstein (Ruge et al., [n. d.]) reorganizes known message sequences to enhance code coverage. He et al. (He et al., 2022) propose a unique fuzzer for the 5G non-access-stratum (NAS) protocol, which extracts packets from captures into a structured message table. This fuzzer then applies different mutation techniques tailored to the types of key fields within the protocol messages, significantly enhancing the intelligence and precision of the message mutation process. For example, length fields are mutated combining boundary and intermediate values, including 0, maximum, minimum, and random intermediate values. It is important to note that mutation-based fuzzers must judiciously manage the correlation of specific fields in message sequences, such as session numbers, counters, or timestamps. Indiscriminate mutations in these fields could render the input ineffective and lead to early rejection. To address this challenge, AFLNET (Pham et al., 2020) modifies the code of the PUT to use a fixed session number 111https://github.com/aflnet/aflnet/blob/master/README.md#step-0-server-and-client-compilation–setup, thus ensuring the effectiveness of the fuzzing process.
@@ 5. Executor
| Year | Work | Tax | Target | Efficient Execution | Runtime Info Extraction | ||||||
@@
| 2020 | Frankenstein(Ruge et al., [n. d.]) | ◐ | Bluetooth | SE | Shared memory | VMSR | - | Code Cov | SDI | ||
@@
| 2020 | ToothPicker(Heinze et al., [n. d.]) | ◐ | Bluetooth | HIL | FHPI | ThrdR | - | Code Cov | SDI | ||
| 2021 | ICS3Fuzzer(Fang et al., 2021) | ● | ICS | - | Socket (P2P) | ProcR | - | Exec Traces | SDI | ||
@@
| 2021 | TCP-Fuzz(Zou et al., [n. d.]) | ◐ | TCP | - | Socket (P2P) | - | - | Branch Cov | SSI | ||
@@
○: Whitebox Fuzzer; ●: Blackbox Fuzzer; ◐: Greybox Fuzzer; Tax: Taxonomy; HIL: Hardware-In-the-Loop; General: The fuzzer is not designed for a specific type of protocol; CUVM: Commonly Used Virtual Machine; SE: Specialized Emulation; OTA: Over-the-air; UDS: Unix Domain Socket; MiTM: Man-in-The-Middle-based packet injection; VMSR: Virtual Machine-level Snapshot and Recovery mechanism; ProcR: Process Restart; PhyR: Physical Reset; ThrdR: Thread Restart; DR: Database Reset; PSR: Process-level Snapshot and Recovery mechanism; UPSR: User-Provided Script Reset; HA: Hardware-Assisted mechanism; EOB: Externally-Observable-Behavior-based method; SDI: Software Dynamic Instrumentation; SSI: Software Static Instrumentation; BAF: Bandwidth Amplification Factor; -: Not implemented; RM: Reset Message; IMFR: In-Memory Filesystem Reset; Resp: Responses.
In this section, we will introduce the key improvements of protocol fuzzers in the executor in detail. As shown in Fig. 6, an executor in protocol fuzzing normally includes four key processes. First, the executor needs to prepare an executable execution environment for PUT (①. Execution Environment Preparation), and then send input to PUT through the input feeding mechanism (②. Input Feeding), extract runtime information during the input processing (③. Information Extraction), and reset the execution state and the environment state to a specific state after the execution of the current iteration is complete (④. Execution Reset).
@@ 5.1. Efficient Execution
In protocol fuzzing, there are commonly two directions to improve the fuzzing efficiency: 1) establishing an execution environment that enables parallel testing for scalability improvement (① in Fig. 6); 2) reducing the execution cost of each testing iteration (② and ④ in Fig. 6).
@@ 5.1.1. Scalability Improvement
Scalable fuzzing, in this context, refers to the capacity to create multiple testing environments for parallel fuzzing. This is crucial in protocol fuzzing, where many fuzzing targets are closely bounded to hardware. The traditional approach for parallel testing of purchasing multiple physical devices can be economically burdensome and inefficient. For protocol fuzzing, since many fuzzing targets depend on specialized execution environments, concurrent testing of these targets can only be carried out by purchasing multiple physical devices, leading to high economic costs and waste.
@@
However, two difficulties hinder the usage of emulation in protocol fuzzing. The first is the availability of the protocol implementation binary, as many firmware images are not publicly available. Second, compared to the diversity of hardware, existing emulators can only support a small fraction of them. These difficulties lead to a lot of work still performing fuzzing in a hardware-in-the-loop way (labeled as “HIL” in Table. 4 column 4) (Luo et al., 2019; Garbelini et al., 2020; Feng et al., 2021; Garbelini et al., 2022c; Gorenc and Molinyawe, 2014; Park et al., 2022).
Some works address these issues according to the characteristics of different devices (labeled as “SE” in Table. 4 column 4). For the first challenge, existing work obtains the target binaries by intercepting Over-The-Air (OTA) firmware updates, or extracting using vendor-specific command or debugging ports. For example, Frankenstein (Ruge et al., [n. d.]) leverages the Patchram mechanism, a Broadcom vendor-specific command that can be used to temporarily patch breakpoints to the ROM, to take the memory snapshot of a physical Bluetooth chip and emulate it in an unmodified version of QEMU. To address the second challenge, the existing work often uses an approach called rehosting to partially emulate the functionality of the physical hardware (Maier et al., 2020; Liu et al., 2021). For example, BaseSafe (Maier et al., 2020) selectively rehosts several signaling message parser functions utilizing the Unicorn engine, which is a popular CPU emulator (Quynh and Vu, 2015).
@@ 5.1.2. Execution Cost Reduction
Another direction to improve fuzzing efficiency is to optimize the intermediate execution steps in each iteration. In the following, we present the progress of existing work towards this direction, which mainly focuses on two subprocedures: input feeding (② in Fig. 6) and execution reset (④ in Fig. 6), specifically.
@@
Input Feeding. Input feeding mechanism acts as a pipeline between the input generator and the PUT to pass the testcase to the PUT for parsing and execution. According to the Inter-Process Communication (IPC) mechanisms that the communication between Fuzzer and PUT relies on, existing approaches can be roughly divided into four categories, namely OTA-based, socket-based, shared-memory-based, and file-based approaches. OTA- and socket-based approaches are mostly used when the fuzzer and PUT cannot be deployed on the same physical device. The latter two approaches can be used to speed up input feeding when the PUT and the fuzzer can be deployed on the same device.
@@
OTA-Based Input Feeding (labeled as “OTA” in Table. 4 column 5). In general, OTA-based input feeding mechanisms are mostly used in the scenario of fuzzing the implementations of protocols that are typically closed in nature and tightly integrated with hardware components, such as Wi-Fi (Garbelini et al., 2022c; Vanhoef, 2017; Sommer et al., 2018), Bluetooth (including classic Bluetooth and BLE) (Garbelini et al., 2022a; Park et al., 2022; Garbelini et al., 2020), LTE (Maier et al., 2020), Zigbee (Ma et al., 2023b), 4G/5G (Garbelini et al., 2022b; He et al., 2022) and SMS/MMS protocols (Gorenc and Molinyawe, 2014). In this approach, PUT and the fuzzer need to be deployed in adjacent physical spaces and communicate with each other on specific frequency bands. Thus, OTA-based fuzzers require the use of radio-frequency transceiver devices with receive and transmit functions, such as a software-defined radio (SDR) to handle signals over a wide tuning range. OTA-based fuzzing provides the capability to test the entire protocol stack including the physical layer. However, OTA-based approaches are the slowest among the above-mentioned approaches. Therefore, many wireless protocol fuzzers try to use other input feeding mechanisms to have better performance, which will be introduced in the following.
@@
Socket-Based Input Feeding (labeled as “Socket” in Table. 4 column 5). Socket-based input feeding mechanisms are mostly used in protocol implementations based on TCP/IP infrastructures. In common cases under these approaches, the fuzzer and the PUT communicate with each other through IP addresses, via socket mechanisms including TCP socket and UDP socket. The socket-based approaches include two deployment modes, one is point-to-point (P2P) communication between the fuzzer and PUT (Chen et al., 2018; Walz and Sikora, 2020; Fiterau-Brostean et al., 2020; Pham et al., 2020; Natella, 2022; Zou et al., [n. d.]; Meng et al., 2021; Li et al., 2022a; Andronidis and Cadar, 2022). The fuzzer can play the role of a client or server depending on the role of the PUT. The other deployment mode is Man-in-the-middle (MiTM), in which the fuzzer acts as a proxy between two communication parties and performs mutation or injection to the normal communication traffic (Jero et al., 2018; Peterson et al., 2020). The MiTM-based input feeding is mostly used in the scenario where the protocol involves certain contextual information (checksum, packet sequences, etc.) that cannot keep valid by mutating static seeds. However, both modes need to address two challenges. First, socket communication is quite heavy and involves lots of context switches. Existing works improve the efficiency of the socket-based input feeding mechanisms by avoiding the use of these expensive network functions. For example, SnapFuzz (Andronidis and Cadar, 2022) replaces the original internet socket with UNIX domain socket (Coffield and Shepherd, 1987), a lightweight IPC mechanism that does not have the routing, checksum calculation operations that IP sockets have. Second, it’s hard for the fuzzer to determine whether the PUT has already finished processing the previous message and is ready to receive the next message. The PUT may reject the messages coming too early when the target is not ready, thus causing the fuzzer to desynchronize from its state machine. To solve this issue, Fiterau-Brostean et al. (Fiterau-Brostean et al., 2020) and AFLNET (Pham et al., 2020) set static time intervals to wait for the PUT to initialize, process requests, and send responses. However, static timers are too coarse-grained and can waste a lot of time waiting for the timeout, thus slowing down the fuzzing process. SnapFuzz (Andronidis and Cadar, 2022) and AMPFuzz (Krupp et al., 2022) develop a more fine-grained method to inspect the state of the socket. Specifically, they use the function call to related network system calls such as , as a sign of ready to receive the next message. They monitor all these function calls through binary rewriting and compile-time code instrumentation, and then notify the fuzzer to send the next iteration of input.
@@
File-Based Input Feeding (labeled as “File” in Table. 4 column 5). File-based input feeding leverages static or dynamic instrumentation techniques to replace heavy network operations with file operations to achieve a performance boost. For example, Yurong et al. (Chen et al., 2019) transform socket communication to file operations using preloading customized libraries (zardus, 2019) under the circumstance that the source code of the PUT is not available. Similarly, Nyx-net (Schumilo et al., 2022) injects a library into the target to hook the network functions of the target connection to obtain their associated file descriptors and injected fuzzing input to the right place.
@@
Shared-Memory-Based Input Feeding (labeled as “Shared-Memory” in Table. 4 column 5). Shared-Memory-Based input feeding writes the fuzzing input to the address of shared memory and hooks the related functions to read the testcase from shared memory (Maier et al., 2020; Ruge et al., [n. d.]; Fu et al., 2023; Bars et al., 2023). For example, BaseSafe (Maier et al., 2020) executes each generated testcase in a forked copy of the target process, and the input for each run is copied to the appropriate address in the corresponding child process. Similarly, Frankenstein (Ruge et al., [n. d.]) creates a virtual modem to inject custom packets. The fuzzed input is written to the receive buffer in RAM that is mapped to the hardware receive buffer using direct memory access (DMA). Also, HNPFuzzer (Fu et al., 2023) emulates network functions based on shared memory to reduce the time consumption due to message transmission between fuzzer and PUT.
@@
Others. There are also works that rely on specialized communication channels to feed fuzzing inputs. For example, to fuzz the client of the Remote Desktop Protocol (RDP), Park et al. leverage the virtual channel, an abstraction layer in RDP that is used to transport data, to actively send fuzzing input from the server to the client (Park et al., 2019). Song et al. use a media converter to convert traffic between Automotive Ethernet and standard Gigabit Ethernet, and fuzz the SOME/IP protocol stack of the electronic control unit (ECU) (Song et al., 2022), which is a control communication protocol between ECUs.
@@
Based on the analysis of the reset of execution methods of existing work, the process mainly involves three key sub-phases, which are 1. reset time selection, 2. execution state reset and 3. execution environment reset. Firstly, the executor assesses whether the current iteration has been concluded during the reset time selection phase. This determination precedes any actions taken to reset the execution. Once it is confirmed that the current execution cycle is complete, the process continues with the execution state reset and execution environment reset, which reset the runtime state of the PUT and the associated external execution environments, respectively. Below we summarize the progress of existing works in these three key phases separately.